# OmniSVG: A Unified Scalable Vector Graphics Generation Model

### v2

@@ 5 Experiments
Table 3: Quantitative Evaluations. Quantitative results between OmniSVG and current state-of-the-art text-to-SVG and image-to-SVG baseline methods. The bold numbers and underlined numbers represents the best and second best performance repectively. Our OmniSVG model demonstrates superior performance compared SOTA SVG generation baselines. StarVector(8B) fails to generate complex SVGs when evaluated in the complex MMSVG-Character dataset.
Evaluation Dataset Methods Time # Tokens Text-to-SVG Image-to-SVG
@@
Vectorfusion [21] 69s 40k 88.74 0.2798 4.98 0.225
SVGDreamer [59] 260s 100k 75.31 0.2923 5.32 0.256
Chat2SVG [56] 28s 1.7k 70.39 0.3029 5.38 0.256
IconShop [55] 6.8s 3.3k 80.84 0.2359 3.47 0.138
LIVE [33] 170s 18.2k 0.960 0.979 0.034 0.004
DiffVG [27] 21s 19.8k 0.951 0.956 0.056 0.015
GPT-4o [20] 6.9s 0.6k 0.887 0.826 0.179 0.134
StarVector(8B) [41] 45s 3.5k 0.977 0.953 0.048 0.012
OmniSVG(3B) 15s 3.8k 64.83 0.3194 5.49 0.247 0.980 0.954 0.049 0.011
MMSVG-Icon OmniSVG(7B) 23s 4.6k 62.16 0.3278 5.63 0.258 0.984 0.973 0.037 0.008
Vectorfusion [21] 70s 41k 90.03 0.2639 4.67 0.214
SVGDreamer [59] 260s 102k 83.02 0.2795 5.09 0.219
Chat2SVG [56] 35s 2.1k 78.84 0.2891 5.01 0.221
IconShop [55] 7.8s 3.3k 84.04 0.2198 3.21 0.098
LIVE [33] 250s 18.2k 0.959 0.960 0.044 0.011
DiffVG [27] 31s 19.8k 0.929 0.930 0.077 0.021
GPT-4o [20] 8.0s 0.7k 0.801 0.783 0.244 0.199
StarVector(8B) [41] 65s 5.3k 0.807 0.831 0.248 0.181
OmniSVG(3B) 49s 9.7k 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019
MMSVG-Illustration OmniSVG(7B) 63s 10.9k 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Vectorfusion [21] 69s 40k 93.03 0.2218 4.08 0.198
SVGDreamer [59] 260s 100k 86.01 0.2423 4.81 0.205
Chat2SVG [56] 42s 1.7k 82.98 0.2499 4.83 0.209
IconShop [55] 6.8s 3.3k 88.41 0.2001 3.13 0.077
LIVE [33] 190s 18.2k 0.911 0.939 0.109 0.038
DiffVG [27] 21s 19.8k 0.876 0.915 0.163 0.051
GPT-4o [20] 8.1s 0.7k 0.712 0.693 0.365 0.349
StarVector(8B) [41]
OmniSVG(3B) 105s 30.8k 72.18 0.2898 5.19 0.213 0.921 0.917 0.049 0.021
MMSVG-Character OmniSVG(7B) 139s 35k 68.02 0.3088 5.33 0.238 0.932 0.941 0.072 0.008
Figure 4: Qualitative comparison with SOTA methods on Text-to-SVG task. We compare the propose method with SOTA Text-to-SVG methods on our evaluation benchmarks, namely Icon, Illustration and Character. The proposed method outperforms existing state-of-the-art approaches in both instruction-following and the aesthetic quality of the generated SVGs.
Figure 5: Qualitative comparison with SOTA methods on Image-to-SVG task. We compare the propose method with SOTA Image-to-SVG methods on our evaluation benchmarks. Despite generating plausible results on simple icon samples, optimization based methods like DiffVG [27] and LIVE [33] tend to output artifacts on complex images. GPT-4o [20] is only able to generate icon-level SVG even given the complex input image. StarVector [41] is able to generate SVG for the input icon image. However, when inputed the illustration or more complex character image, StarVector fails to generate SVGs. Please zoom-in for more details.
To validate the effectiveness of our method, we first introduce the baselines and implementation details (Sec. 5.1). Then, we make quantitative comparisons with prior arts (Sec. 5.2) and ablation studies (Sec. 5.3) to study the effectiveness of our design.
@@ 5.1 Baselines
For the text-to-SVG task, we compare our method with language-based (LLM-based) methods, including VectorFusion [21], SVGDreamer [59], Chat2SVG [56] and IconShop [55]. For image-to-SVG task, we compare our method with baseline methods across image vectorization and Multimodal Large Language Modeling approaches, including LIVE [33], DiffVG [27], StarVector [41] and GPT-4o [20] using the official implementations with the hyperparameters proposed by the authors, and apply their pre- and post-processing code as required.
Specifically, for the text-to-SVG task, the optimization-based method SVGDreamer excels in enhancing editability by employing a semantic-driven image vectorization process that effectively separates foreground objects from the background, while failing to handle complex scenes. Another optimization-based work, VectorFusion, stands out for generating SVG-exportable vector graphics without relying on large captioned datasets. However, Vectorfusion is also unable to handle complex scenarios and diverse styles. The significant problem with these optimization-based works is that the optimization time is too long. Generating an SVG usually takes more than ten minutes, which is too expensive. For the LLM-based method, Chat2SVG integrates Large Language Models (LLMs) with image diffusion models to create semantically rich SVG templates. However, Chat2SVG still needs to optimize the output SVG script from LLM, which introduces increased computational complexity and poses challenges during model training. In comparison, IconShop utilizes a transformer-based architecture to autoregressively model SVG path sequences, demonstrating exceptional performance in simplified icon SVGs, which offers effective solutions for text-to-SVG generation. It can only generate black simple Icon SVGs.
For the image-to-SVG task, we compare our method with the image vectorization methods. LIVE allows progressive and efficient generation of SVGs, optimizing closed vector paths under raster image supervision with shape complexity control. However, LIVE needs to optimize for a long time when generating complex SVGs. DiffVG enables end-to-end differentiability in vector graphics rasterization, improving optimization through anti-aliasing and gradient-based methods while also is computationally expensive due to the complexity of the forward-backward rasterization process. Recently, the Multimodal Large Language Model (MLLM) based method StarVector leverages the visual understanding to apply accurate SVG primitive to the LLM architecture, which also can generate SVGs from both text and image inputs. However, it still fails to generate complex SVGs. Since Starvector [41] has not yet opened up its text-to-SVG model weights, our MMSVGBench does not evaluate Starvector’s text-to-SVG capabilities. MMSVG-Bench also evaluates our methods with VLM methods, GPT-4o, to conduct a comprehensive assessment. We compare our method with these baselines on our MMSVG-2M dataset, from simple MMSVG-Icon datset, a bit complex MMSVG-illustration dataset, to the very complex MMSVG-Character dataset. We report the quantitative results in Tab. 3 and the qualitative results in Fig. 4 and Fig. 5.
@@
Figure 6: Generated SVG with Character-Reference (CRef) by OmniSVG. By training on MMSVG-Character with natural character image and SVG pair data, OmniSVG is capable of generating character SVGs through image references.
5.2 Evaluations and Comparisons
5.2.1 Quantitative Evaluations
We report the evaluation metrics of our OmniSVG and baselines on our MMSVG-2M dataset, including the MMSVG-Icon, MMSVG-Illustration and MMSVG-Character in Tab. 3, including both text-to-SVG and image-to-SVG tasks. As shown in Tab. 3, OmniSVG consistently outperforms all other text-to-SVG baselines across all evaluation metrics, indicating lower FID values, higher CLIP scores, higher aesthetic scores and higher HPS scores. For the image-to-SVG tasks, while image vectorization method LIVE achieves higher SSIM, lower LPIPS and lower MSE scores, our OmniSVG can achieve higher DINO scores, indicating the higher semantic features between the input image and the generated SVGs.
5.2.2 Qualitative Evaluations
Text-to-SVG task As shown in Fig. 5, we compare the performance of our method with several baseline approaches using seven distinct text prompt instructions for the text-to-SVG task. For optimization-based methods, such as SVGDreamer [59] and VectorFusion [21], generating an SVG from a single prompt requires significant computation time. These methods rely on iterative optimization processes, which, while effective for refining SVG details, are computationally expensive and time-consuming. In contrast, language-based methods like IconShop [55] and Chat2SVG [56] offer faster generation times, as they leverage pre-trained models to generate SVG templates more efficiently. However, each of these methods comes with notable limitations. IconShop, despite producing complete SVG shapes, is restricted to generating black-and-white graphics, which limits its applicability in scenarios requiring color and visual richness. Chat2SVG, although more flexible, is a two-stage process. In this section, we focus on the first stage of Chat2SVG, which is responsible for generating an initial SVG template. While it offers some flexibility, the generated SVGs often lack the level of detail and semantic consistency observed in the outputs from our method.
Our method, OmniSVG, consistently outperforms all other baselines across various text prompts. The generated SVGs not only retain high fidelity to the input instructions but also incorporate rich color and geometric accuracy. Additionally, OmniSVG demonstrates the ability to handle more complex visual cues, enabling the generation of both simple and intricate SVG designs. This qualitative evaluation highlights the superior performance of our approach in terms of both speed and the quality of the resulting SVGs, making it a more effective and versatile solution compared to existing methods.
Image-to-SVG Task. For the image-to-SVG task, we compare our method with classical image vectorization approaches, including DiffVG [27], LIVE [33], VLM-based methods GPT-4o [20] and StarVector [41]. As shown in Fig. 5, our method outperforms these baselines in both the quality and efficiency of SVG generation.
We present a comprehensive qualitative comparison of our proposed method against state-of-the-art (SOTA) Image-to-SVG techniques, evaluated on our established benchmark datasets. The methods under consideration include optimization-based approaches such as DiffVG [27] and LIVE [33], as well as neural-based methods like GPT-4o [20] and StarVector [41]. While image vectorization methods such as DiffVG and LIVE demonstrate impressive results when applied to simple icon images, they exhibit notable deficiencies when tasked with more complex images. These methods tend to generate visual artifacts, indicating that they struggle with preserving the integrity of intricate visual structures, particularly when the input images feature detailed shapes, gradients, or textures. The GPT-4o model, though capable of generating SVG outputs for complex input images, is constrained to producing only icon-level SVGs. This limitation suggests that GPT-4o is primarily designed to handle simpler iconography and lacks the capacity to fully interpret and convert detailed illustrations into scalable vector graphics, which may be a result of its underlying design or training data constraints. StarVector, another VLM-based method, demonstrates strong performance when tasked with generating SVGs from simple icon images. However, its effectiveness diminishes significantly when dealing with more complex inputs, such as illustrations or intricate character images. In these cases, StarVector fails to generate accurate or meaningful SVG outputs, highlighting its limited generalization ability for more diverse and intricate image types.
While our OmniSVG are able to efficiently convert images into high-quality, editable SVGs, from the input icon images, illustration images to the complex character images, highlighting the superior performance in handling diverse and intricate visual cues, setting it apart from traditional vectorization approaches. We provide more visual results in Fig. 8.
Figure 7: Qualitative study on parametrization. Ablation studies on color parametrization (abbreviated as param.) and coordinate (abbreviated ad coord.) paramterization are conducted.
Character-Reference SVG generation task. As shown in Fig. 6, by training on MMSVG-Character with natural character image and SVG pair data, OmniSVG is capable of generating character SVGs through image references. Moreover, we assess the alignment between input character images and generated SVGs using the third-party VLMs assistant GPT-4o [20]. Our findings reveal an average score of approximately $7$, suggesting that the generated IP SVGs exhibit a well degree of similarity to the input reference images.
@@ 5.3 Ablation studies
w/o param. 97.01 0.1537 2.873 0.112 0.702 0.668 0.334 0.232 35k
w/o coordinate param. 87.93 0.2193 3.352 0.156 0.783 0.769 0.295 0.160 29.5k
w/o color param. 76.39 0.2632 4.234 0.201 0.901 0.912 0.098 0.028 13.1k
full param. (Ours) 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019 9.7k
We conduct ablation studies on the SVG parameterization, the scaling up model size and the VLM architecture to validate the effectivenss of our design. Most of our ablation experiments are performed on the OmniSVG-3B model for faster iteration, and we empirically find they work well on the larger OmniSVG-7B.
Effectiveness of the SVG Parameterization. In this experimental study, we present a comprehensive comparison between SVG parameterization approach and traditional non-parameterized methods for SVG representation in large language models. The fundamental distinction lies in how coordinate information is tokenized: traditional methods represent each digit as a separate token (e.g., coordinates $(123,456)$ require six tokens plus delimiters), whereas our parameterization method encodes entire coordinates as single semantic tokens, dramatically improving the model’s ability to process complex path information.
We parameterize both the coordinate and color attributes of the SVG. To validate the effectiveness of our experimental design, we conducted the following ablation experiments: parameterizing color while leaving the coordinates unparameterized, parameterizing coordinates while leaving the color unparameterized, not parameterizing either the coordinates or the color, and parameterizing both the coordinates and the color (our method). The results of these experiments on the MMSVG-Illustration dataset are summarized in the Tab. 4 and Fig. 7.
The results in Tab. 4 indicate that incorporating both coordinate and color parameterizations yields the most favorable outcomes across all evaluated metrics. Notably, this configuration achieves the lowest token length, suggesting that our method efficiently utilizes token space. This efficiency implies that our approach can generate complex SVGs with reduced computational resources, which is advantageous for applications requiring scalability and efficiency. Moreover, the qualitative results in Fig. 7 indicates that the both coordinate and color parameterizations significantly outperforms other methods, particularly as SVG complexity increases. No parameterizations method fail to generate SVGs when given the complex image. These experiments underscore the effectiveness of our full parameterization strategy in balancing performance and resource utilization, making it a compelling choice for SVG generation tasks.
Ablation studies on the model size. To analyze whether training a larger model benefits SVG generation, we evaluate OmniSVG base models with different sizes on the MMSVG-2M dataset in Tab. 5. We evaluate OmniSVG with base models of varying sizes on the MMSVG-2M dataset in Tab. 5. We progressively scale up the model size from the FLAN-T5-base [10] (223M) to Qwen2.5-VL-3B-Instruct (3.7B) and Qwen2.5-VL-7B-Instruct (8.3B). We observe that as the model size grows, the quality of the generated samples gets higher, indicating the effectiveness of scaling up generative auto-regressive training on multi-modal SVG generation tasks.
@@
FLAN-T5-Base[10] Text 223M 89.30 0.177 3.437 0.102
FLAN-T5-Large[10] Text 770M 83.92 0.2303 4.102 0.177
FLAN-T5-xl[10] Text 3B 76.92 0.2789 4.642 0.217
blip2-flan-t5-xl[26] Text/Image 3.94B 75.23 0.2881 4.712 0.201 0.891 0.902 0.099 0.046
OmniSVG(3B) Text/Image 3.7B 70.45 0.3077 5.395 0.245 0.974 0.944 0.069 0.019
OmniSVG(7B) Text/Image 8.3B 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Ablation studies on the VLM architecture. To illustrate the effectiveness of the VLM architecture, we conducte an ablation study by replacing the VLM architecture with alternative LLM-based architectures, incorporating image encoders such as CLIP ViT-B/32 [37], VQGAN [14], and the Qwen2.5-VL [1] model used in OmniSVG. These experiments were carried out on the MMSVG-Illustration dataset, with a focus on comparing the performance of different architectures in generating SVG outputs. Among these architectures, Qwen2.5-VL consistently delivered superior results across all evaluation metrics.
The results in Tab. 6 clearly indicate that none of the alternative architectures outperformed Qwen2.5-VL, highlighting the importance of selecting an architecture with a robust image encoder. Additionally, a key advantage of Qwen2.5-VL is its support for multi-image input, which enables the generation of more complex outputs, such as character reference SVGs.
Table 6: Ablation of the VLM architecture. we ablate the VLM architecture by replacing it with the LLM architecture with the image encoder, including the CLIP ViT-B/32 [37], VQGAN [14], and the Qwen2.5-VL [1] we use in OmniSVG. We experiment with training experiments on the MMSVG-Illustration dataset. Qwen2.5-VL architecture gives the best results on all metrics.
Vision Language Text-to-SVG Image-to-SVG
@@
CLIP Qwen2.5 72.243 0.2876 4.57 0.209 0.881 0.872 0.176 0.062
VQGAN Qwen2.5 76.29 0.2739 4.62 0.198 0.890 0.852 0.183 0.065
Qwen2.5-VL-3B-Instruct 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019
Qwen2.5-VL-7B-Instruct 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Table 7: User Study of OmniSVG and baselines. Our OmniSVG achieves the user preference among the baselines.
@@
Vectorfusion [21] 40 60 80
SVGDreamer [59] 40 70 80
Chat2SVG [56] 60 60 90
IconShop [55] 80 60 80
GPT-4o [20] 40 60 80
StarVector(8B) [41] 40 80 70
DiffVG [27] 90 80 100
LIVE [33] 90 70 100
OmniSVG [33] 100 90 100
5.4 User study and Implementation Details.
User Study. For user study, we assess the user preference, vividity, and text-SVG/image-SVG alignment of the generated SVGs. In this study, we conduct a user study to evaluate the effectiveness and practicality of the SVGs generated by our method and baselines. Specifically, We randomly sample 5 text descriptions and 3 image prompts from the evaluation dataset and generate corresponding SVGs using our methods and baseline models. Along with the original 8 SVGs from the dataset, this provided a total of 64 SVGs for the user study. In Tab. 7, preference indicates the generated SVG preference, “Vividlity” indicates the sample is more lively, and “Alignment” represents the input text/image instructions more likely to align with the generated SVGs. In this study, we shuffle the SVGs generated by the different models together with the selected examples from our dataset. Participants are only provided with the text descriptions or image prompts and the shuffled SVGs. Tab. 7 also demonstrates that OmniSVG is preferred by volunteers, who perceive the generated SVGs as more vivid and visually aligned with the input text or image. This alignment highlights OmniSVG’s superior ability to produce more accurate and contextually relevant SVG content according to volunteers’ feedback.
Implementation Details. We train our models in bfloat16 with the ZeRO-2 strategy [39] for memory-efficient training. We also adopt the AdamW [32] optimizer with a learning rate decaying from $3\times 10^{-4}$ to $3\times 10^{-6}$ and a weight decay of $0.1$ to train our model. In practice, we load the pre-trained weights from the Qwen2.5-VL [1] model and initialize the SVG embeddings from scratch. Without further specification, we generate SVGs with the top-k and top-p sampling strategy with $k=50$ and $p=0.95$ for diversity.

### v3

@@ 5 Experiments
Table 2: Quantitative Evaluations. Quantitative results between OmniSVG and current state-of-the-art text-to-SVG and image-to-SVG baseline methods. The bold numbers and underlined numbers represents the best and second best performance repectively. Our OmniSVG model demonstrates superior performance compared SOTA SVG generation baselines.
Evaluation Dataset Methods # Tokens Text-to-SVG Image-to-SVG
@@
Vectorfusion vectorfusion 66.2k 250.77 0.240 4.76 0.237
SVGDreamer svgdreamer 132.0k 308.94 0.207 4.26 0.221
Chat2SVG chat2svg 0.6k 190.87 0.299 4.41 0.247
IconShop iconshop 2.0k 213.28 0.288 4.55 0.244
LIVE live 52.5k 0.932 0.943 0.106 0.011
DiffVG diffvg 322.0k 0.940 0.954 0.066 0.002
GPT-4o GPT-4o 0.3k 0.860 0.792 0.403 0.124
StarVector(8B) starvector 2.0k 0.895 0.881 0.231 0.059
Vtracer 52.4k 0.993 0.966 0.039 0.002
OmniSVG(4B) 3.8k 137.40 0.275 4.62 0.244 0.993 0.950 0.050 0.006
MMSVG-Icon OmniSVG-L(8B) 5.7k 130.56 0.276 4.60 0.242 0.922 0.893 0.235 0.040
Vectorfusion vectorfusion 66.1k 253.94 0.185 4.94 0.226
SVGDreamer svgdreamer 132.0k 419.70 0.201 4.37 0.221
Chat2SVG chat2svg 1.0k 210.03 0.283 4.45 0.250
IconShop iconshop 2.6k 107.93 0.233 4.46 0.224
LIVE live 52.2k 0.935 0.950 0.111 0.008
DiffVG diffvg 322.0k 0.945 0.955 0.065 0.001
GPT-4o GPT-4o 0.4k 0.875 0.854 0.373 0.077
StarVector(8B) starvector 2.6k 0.877 0.900 0.238 0.046
Vtracer 57.6k 0.994 0.966 0.035 0.002
OmniSVG(4B) 5.8k 154.37 0.226 4.56 0.232 0.899 0.906 0.237 0.034
MMSVG-Illustration OmniSVG-L(8B) 6.9k 138.42 0.231 4.51 0.232 0.905 0.907 0.231 0.031
To validate the effectiveness of our method, we first introduce the baselines (Sec. 5.1). Then, we make quantitative comparisons with prior arts (Secs.˜5.3 and 5.2) and conduct ablations (Sec. 5.4) to study the effectiveness of our design.
@@ 5.1 Baselines
For the text-to-SVG task, we compare our method with language-based (LLM-based) methods, including VectorFusion (vectorfusion), SVGDreamer (svgdreamer), Chat2SVG (chat2svg) and IconShop (iconshop). For image-to-SVG task, we compare our method with baseline methods across image vectorization and Multimodal Large Language Modeling approaches, including LIVE (live), DiffVG (diffvg), StarVector (starvector), Vtracer (vtracer) and GPT-4o (GPT-4o) using the official implementations with the hyperparameters proposed by the authors, and apply their pre- and post-processing code as required.
5.2 Quantitative Comparisons
We compare our OmniSVG with other baseline methods on the “text-to-SVG” and “image-to-SVG” tasks in our MMSVG-Bench. In addition to the metrics mentioned in Sec.˜3, we also report the average token length (# tokens) of a generated SVG sample utilizing the Qwen2.5-VL (qwen2.5-VL) tokenizer.
As shown in Tab. 2, OmniSVG demonstrates strong performance compared to state-of-the-art baselines in text-to-SVG generation, achieving superior FID scores and competitive CLIP score, aesthetic quality, and HPS. For image-to-SVG, OmniSVG also achieves competitive results with traditional vectorization methods, i.e. LIVE live, DiffVG diffvg, and VTracer vtracer, but with a much shorter sequence length. When comparing to auto-regressive methods, i.e. GPT-4o GPT-4o and StarVector starvector, OmniSVG also achieves a superior performance across all metrics. The above results indicate that OmniSVG effectively balances the generation cost and the visual quality when generating SVGs according to multi-modal conditions.
5.3 Qualitative Evaluations
Text-to-SVG task
. We compare our method with baseline approaches using seven distinct text prompts for the text-to-SVG task, as shown in Fig.˜4. Optimization-based methods like SVGDreamer (svgdreamer) and VectorFusion (vectorfusion) require significant computation time due to their iterative optimization processes, which, while effective for refining SVG details, are computationally expensive. Auto-regressive methods, such as IconShop (iconshop) and Chat2SVG (chat2svg), generate SVGs more quickly by leveraging pre-trained models but have notable limitations. IconShop produces monochrome SVGs, restricting its applicability, while Chat2SVG, though flexible, generates less detailed and semantically consistent SVGs in its first stage. Our OmniSVG consistently outperforms all baselines across various text prompts in generating high-fidelity SVGs with rich color, geometric accuracy, and the ability to handle complex visual cues.
Figure 3: Qualitative Comparison with SOTA Methods on Text-to-SVG Task. We compare the propose method with SOTA Text-to-SVG methods on our evaluation benchmarks, namely Icon and Illustration.
Figure 4: Qualitative Comparison with SOTA Methods on Image-to-SVG Task. We compare the propose method with SOTA Image-to-SVG methods on our evaluation benchmarks.
Figure 5: Generated SVG with Character-Reference (CRef) by OmniSVG.
Image-to-SVG Task.
We compare our method with classical image vectorization approaches, including DiffVG (diffvg), LIVE (live), and VLM-based methods GPT-4o (GPT-4o), StarVector (starvector) and Vtracer (vtracer) As shown in Fig.˜4, our method outperforms these baselines in both quality and efficiency. Optimization-based methods like DiffVG and LIVE perform well on simple icons but struggle with complex images, often generating visual artifacts. The GPT-4o model, while capable of generating SVGs for complex images, is limited to icon-level outputs and cannot handle detailed illustrations. StarVector excels at simple icons but fails to produce accurate SVGs for more intricate images, highlighting its limited generalization capability. Vtracer is an image processing algorithm designed to convert raster images into SVGs. In contrast, OmniSVG efficiently converts a wide range of images, from icons to complex illustrations and character images, into high-quality, editable SVGs. This superior performance in handling diverse visual cues distinguishes OmniSVG from traditional vectorization methods. Additional visual results can be found in Fig.˜12. We provide more detailed discussions with existing methods, particularly the recent works LLM4SVG llm4svg and StarVector starvector, in the Appendix˜D.
Character-Reference SVG generation task.
@@ 5.3 Qualitative Evaluations
As shown in Fig.˜5, by training on MMSVG-Character with natural character image and SVG pair data, OmniSVG is capable of generating character SVGs through image references.
@@ 5.4 Ablation studies
w/o param. 218.76 0.185 3.43 0.138 0.741 0.718 0.315 0.182 18.5k
w/o coordinate param. 193.42 0.216 3.90 0.169 0.826 0.809 0.248 0.119 10.2k
w/o color param. 167.28 0.269 4.31 0.211 0.895 0.879 0.179 0.053 6.3k
OmniSVG(4B) 145.89 0.308 4.59 0.238 0.946 0.928 0.138 0.020 4.8k
Effectiveness of SVG Parameterization. We present a comprehensive comparison among different SVG parameterization strategy with the traditional non-parameterized methods for SVG representation in large language models. We ablates on the parameterization on both coordinate and color attributes of the SVG.
The results, shown in Tab.˜3 and Fig.˜6 demonstrate that parameterizing both coordinate and color attributes yields a better generation results under all metrics with the shortest token length. It further validates that the efficient token representation allows our method to generate complex SVGs with fewer computational resources. Additionally, qualitative results show that our method outperforms others, particularly as SVG complexity increases. The non-parameterization method fails to generate SVGs for complex images. These findings underscore the effectiveness of our full parameterization strategy in balancing performance and resource efficiency for SVG generation tasks.
@@
FLAN-T5-Base(t5) Text 223M 198.48 0.158 3.38 0.085
FLAN-T5-Large(t5) Text 770M 175.24 0.208 3.92 0.142
FLAN-T5-xl(t5) Text 3B 160.28 0.258 4.31 0.192
blip2-flan-t5-xl(blip2) Text/Image 3.94B 152.11 0.235 4.48 0.215 0.898 0.891 0.255 0.041
OmniSVG(4B) Text/Image 3.7B 145.89 0.308 4.59 0.238 0.946 0.928 0.138 0.020
Ablation studies on model size. To analyze whether training a larger model benefits SVG generation, we evaluate OmniSVG base models with different sizes on the MMSVG-2M dataset in Tab.˜4. We evaluate OmniSVG with base models of varying sizes on the MMSVG-2M dataset in Tab. 4 by progressively scaling up the model size. The results show that as the model size grows, we can generate SVG samples with a better quality.
Table 5: Ablation on VLM architecture.
Vision Model Language Model Text-to-SVG Image-to-SVG
@@
CLIP Qwen2.5 185.31 0.249 4.52 0.215 0.867 0.856 0.267 0.058
VQGAN Qwen2.5 198.74 0.234 4.49 0.203 0.839 0.828 0.295 0.071
Qwen2.5-VL-3B-Instruct 145.89 0.308 4.59 0.238 0.946 0.928 0.138 0.020
Qwen2.5-VL-7B-Instruct 134.45 0.254 4.56 0.237 0.914 0.900 0.233 0.036
Ablation Studies on the VLM Architecture. To evaluate the effectiveness of the VLM architecture, we conducted an ablation study replacing it with alternative LLM-based architectures incorporating image encoders such as CLIP ViT-B/32 (clip-vit/32), VQGAN (vqgan), and Qwen2.5-VL (qwen2.5-VL).
Figure 6: Qualitative Study on Parametrization.
The results in Tab.˜5 show that Qwen2.5-VL consistently outperformed all alternatives under all evaluation metrics.
Figure 7: User Study of OmniSVG and baselines.
@@
Vectorfusion vectorfusion 35 58 76
SVGDreamer svgdreamer 41 65 79
Chat2SVG chat2svg 55 61 86
IconShop iconshop 79 57 75
GPT-4o GPT-4o 38 54 80
StarVector(8B) starvector 37 81 68
DiffVG diffvg 88 76 96
LIVE live 86 70 95
OmniSVG 96 88 98
User Study. We extract one-tenth of the samples from the evaluation dataset and conducted a user study with 15 participants to evaluate user preferences, vividness, and the alignment between text-to-SVG and image-to-SVG. Participants are asked to assess SVGs generated by different models based on 150 text descriptions and 150 image prompts, comparing the results generated using our method and baseline models. The results in Fig.˜7 show that OmniSVG is widely preferred, with higher scores for vividness and superior semantic alignment with the input conditions.